Calculate the concentration of sodium chloride present in mass % for a solution of 45 g of salt in 1200 mL of water.

The density of water is 1.00 g/mL.


1
Expert's answer
2021-08-04T07:13:29-0400

Moles of sodium s chloride ="\frac{45}{58.44}=0.77moles"


Volume "=\frac{1200}{1000}=1.2L"


Molarity "=\frac{0.77}{1.2}=0.64M"